OpenStudy (thebeatles1200):

Multiplying Mixed numbers! Please help! 4 1/3 times 2 1/4? I got 29 1/4 as the answer when the correct answer is suppose to be 9 3/4 or 39/4.

OpenStudy (thebeatles1200):

So I first convert the mixed numbers to improper fractions and I get 13/1 and 9/4 so I then multiply and get 117/4 which then equals 29 1/4 as my answer. What am I doing wrong?

OpenStudy (mathmale):

Just great that you have remembered and used material from previous problems!! Check your first "improper fraction;" you have an arithmetic mistake there.

OpenStudy (mathmale):

Convert 2 1/4 to an improper fraction.

OpenStudy (thebeatles1200):

2 times 4 is 8 plus 1 equals 9/4?

OpenStudy (thebeatles1200):

Did I covert it wrong?

OpenStudy (mathmale):

that's better. What are your 2 improper fractions? Show the product of these 2 improper fractions.

OpenStudy (thebeatles1200):

13/1 times 9/4 equals 117/4

OpenStudy (mathmale):

Like this: (fraction 1) * (fraction 2)

OpenStudy (thebeatles1200):

13/1 * 9/4 equals 117/4

OpenStudy (mathmale):

But one of the mixed numbers is 4 1/3. What happened to the denominator, 3?

OpenStudy (mathmale):

\[\frac{ 13 }{ ? }*\frac{ 9 }{ 4 }=?\]

OpenStudy (thebeatles1200):

But we turned that into a improper fraction so the denominator became 1...

OpenStudy (thebeatles1200):

13/1

OpenStudy (mathmale):

Not so.

OpenStudy (mathmale):

Try again.

OpenStudy (thebeatles1200):

Umm is there something I need to do after converting them to improper fractions to make them have the same denominator?

OpenStudy (mathmale):

Here's what we do with 4 1/3: 1. Multiply that 4 by 3. 2. Add 1 to your result. 3. Write this latest result as the numerator of a fraction with 3 in the denom.

OpenStudy (mathmale):

One question at a time. Please convert 4 1/3 to an improper fraction.

OpenStudy (thebeatles1200):

Ohhh okay I see what I did wrong with that

OpenStudy (mathmale):

What are the 2 improper fractions that you're multiplying?

OpenStudy (thebeatles1200):

13/3 and 9/4

OpenStudy (thebeatles1200):

13/3 *9/4 = 117/12 right?

OpenStudy (mathmale):

Right. 13/3 is correct/ 13/1 is not. Now write the product of 13/3 and 9/4.

OpenStudy (thebeatles1200):

117/12 but that has to be simplified right?

OpenStudy (mathmale):

Try this: \[\frac{ 13 }{ 3 }*\frac{ 9 }{ 4 }\]

OpenStudy (mathmale):

Instead of multiplying 13 by 9, CANCEL the 9 and the 3. Be careful.

OpenStudy (thebeatles1200):

Wait, cancel the 9 and 3?

OpenStudy (mathmale):

3*3/3 = ?

OpenStudy (thebeatles1200):

Dont we just multiply 13 times 9 and 3 times 4

OpenStudy (thebeatles1200):

3

OpenStudy (mathmale):

\[\frac{ 3(3) }{ 3 }=?\]

OpenStudy (thebeatles1200):

9/3=3

OpenStudy (mathmale):

What is your final result? 13/3 times 9/4

OpenStudy (thebeatles1200):

39/4

OpenStudy (mathmale):

That's it. So much easier to cancel instead of multiplying.

OpenStudy (thebeatles1200):

Thanks so much for helping me find my error!

OpenStudy (mathmale):

My pleasure. Keep up the good work and the good attitude.

OpenStudy (thebeatles1200):

Thank you! Have a great evening!